Grade 12th passMechanics

a man throws balls with the same speed vertically upwards one after the other at an interval of 2 seconds. what should be the speed of the throw so that more than two balls are in the sky at any time?

Jagadish raj
12 Years agoGrade 12th pass

1 Answer

Sumit Majumdar
12 Years ago
Dear student,
If we are to say that there should be atleast two balls in air, then, we can say that one is just thwon and the other is at the highest point. Then the time interval between them is 2 sec.

Hence using:
v=u-gt\Rightarrow 0=u-gt\Rightarrow u=gt=9.8\times 2=19.6 m/s
Hence for more than 2 balls in air, the speed should be more than 19.6 m/s,
Regards
Sumit